# Non-Stationary Lipschitz Bandits

### v1

@@ D.6 Upper bounding (A)
Our approach begins by upper bounding (A) through a decomposition at depth $d_{0}(t)$ and leverage the properties of our sampling scheme (Algorithm 2) . In particular, it leverages the fact that at each round $t$, we choose uniformly an active bin at depth $d_{0}(t)$.
@@
where in the last inequality we used the fact that $B_{t,m}\subseteq\mathrm{Parent}(B_{t,m},d_{0}(t))$, $B^{\rm{last}}_{l,m}\subseteq\mathrm{Parent}(B^{\rm{last}}_{l,m},d_{0}(t))$ and Assumption 1. Note that $\mathrm{Parent}(B_{t,m},d_{0}(t))$ is necessarily active at depth $d_{0}(t)$. Summing this inequality over the block $[\tau_{l,m},\tau_{l,m+1}[$,
@@
We are now ready to prove Proposition 7. The rest of this subsection is dedicated to this purpose.
Thanks to Lemma 2, it now suffices to bound both the parent’s regret contribution at depth $d_{0}(t)$
@@
Upper bounding (A.1.). Considering all possible value of $d_{0}(t)$, (A.1) can be rewritten using our ordering introduced in Definition 5: denoting $B_{d}^{(1)}$, $B_{d}^{(1)}\dots,B_{d}^{(2^{d})}$ the bins at depth $d$,
[…]
@@
For all bin $B$, $M(s,d,B)$ denotes the last round where bin $B$ is active within a replay at depth $d$ starting at round $s$, $M(s,d,B)\in[s,s+8^{d}]$. If bin $B$ is not evicted during this replay, we define $M(s,d,B)=s+8^{d}$. By abuse of notations, we define $M(\tau_{l,m},m,B)$ as the last round where bin $B$ is retained in $\mathcal{B}_{\mathrm{MASTER}}(m)$, and define $M(\tau_{l,m},m,B)=\tau_{l,m}+8^{m}-1=\tau_{l,m+1}-1$ if it is not evicted during the block $[\tau_{l,m},\tau_{l,m+1}[$.
@@
Without any loss of generality, we assume that $\{B_{m}^{(1)},\cdots,B_{m}^{(2^{m})}\}$ is the ordering according to the eviction time of these bins in the block $[\tau_{l,m},\tau_{l,m+1}[$ (otherwise we can always re-index the bins), that is,
@@
Moreover, since $B^{\rm{last}}_{l,m}$ and $B_{m}^{(k)}$ are by definition not evicted on interval $[\tau_{l,m},M(\tau_{l,m},m,B_{m}^{(k)})]$, we have by Proposition 5 on event $\mathcal{E}_{1}$
@@
$\leq(c_{0}+c_{1}+8)\log(T)4^{m}\,,$
assuming again horizon $T$ satisfies $\log(T)\geq 1$. Therefore, the regret contribution of replay at depth $m$ is upper bounded as
$\sum_{k=1}^{2^{m}}\sum_{t=\tau_{l,m}}^{M(\tau_{l,m},m,B_{m}^{(k)})}\mathds{1}\!\left\{d_{0}(t)=m\right\}\frac{\delta_{t}\left(B^{\rm{last}}_{l,m},B_{m}^{(k)}\right)}{|\mathcal{B}_{t}(m)|}$ $\leq(c_{0}+c_{1}+8)\log(T)4^{m}\sum_{k=1}^{2^{m}}\frac{1}{2^{m}+1-k}$
@@
$\leq(c_{0}+c_{1}+9)\log^{2}(T)4^{m}\,.$ (14)
Regret contribution of replays at depth $d<m$. Fix a depth $d\in[\![0,m-1]\!]$ and a bin $B_{d}^{(k)}$ at this depth. We consider the set of rounds $t>\tau_{l,m}$ for which $d_{0}(t)=d$ and $B_{d}^{(k)}\in\mathcal{B}_{t}(d)$, and partition this set into contiguous intervals (represented as blue intervals in Figure 1).
@@
Each interval concludes either when $B_{d}^{(k)}$ is evicted from the active set $\mathcal{B}_{t}(d)$, or when the replay itself ends at time $s+8^{d}$. Therefore, each interval spans the range $[s,M(s,d,B_{d}^{(k)})]$, where $M(s,d,B_{d}^{(k)})$ is the final round during which $B_{d}^{(k)}$ is active within that specific replay (see Definition 11).
@@
Following these remarks, we can upper bound the second sum of Section D.6 on event $\mathcal{E}_{1}$ as
@@
Now, since the bin $B_{d}^{(k)}$ is active over the interval $[s,M(s,d,B_{d}^{(k)})]$, we can apply Proposition 5 to obtain the desired bound on the event $\mathcal{E}_{1}$.
@@
$\leq(c_{0}+c_{1}+8)\log(T)4^{d}\,,$
assuming horizon $T$ satisfies $\log(T)\geq 1$, and where the second inequality follows from the fact that the length of a replay at depth $d$ stating at round $s$ is at most $8^{d}$. Therefore, we have
@@
$\leq(c_{0}+c_{1}+8)\log(T)\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{d=0}^{m-1}\sum_{k=1}^{2^{d}}\frac{1}{2^{d}-k+1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}4^{d}\,\bigg{|}\,\mathcal{F}_{\tau_{l,m}}\right]$
$\leq(c_{0}+c_{1}+8)\log(T)\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{d=0}^{m-1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}(\log(2^{d})+1)4^{d}\,\bigg{|}\,\mathcal{F}_{\tau_{l,m}}\right]$
$\leq(c_{0}+c_{1}+8)\log^{2}(T)\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{d=0}^{m-1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}4^{d}\,\bigg{|}\,\mathcal{F}_{\tau_{l,m}}\right]$
$\leq(c_{0}+c_{1}+8)\log^{2}(T)\mathbb{E}\left[\sum_{d=0}^{m-1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}4^{d}\,\bigg{|}\,\mathcal{F}_{\tau_{l,m}}\right]\,,$
@@
We now leverage the fact that relatively few replays occur in expectation, and the fact that the Bernoulli random variables $R_{s,d}$ are sampled independently of the observations collected during the episode. More precisely, all Bernoulli variables are drawn i.i.d. at the beginning of the block, i.e. at round $t=\tau_{l,m}$. This allows us to treat the scheduling decisions as fixed ahead of time. Furthermore, by the design of the replay scheduling mechanism in Algorithm 2,
@@
$\leq 6\sqrt{8}(c_{0}+c_{1}+8)\log^{2}(T)4^{m}\,.$ (16)
From Section D.6, combining Sections D.6 and D.6 gives the desired rate for (A.1),
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(A.1)}\leq 18(c_{0}+c_{1}+9)\log^{2}(T)4^{m}\,.$ (17)
Upper bounding (A.2). To bound this term, we apply the same reasoning previously used to control (A.1), and in particular the partition argument of Figure 1. By observing that each replay at depth $d_{0}(t)=d$ is initiated by a round $s$ such that $R_{s,d}=1$, and by using the fact that each replay at depth $d$ lasts at most $8^{d}$ rounds, we distinguish the cumulative discretization bias incurred at depth $m$ and the one incurred at other depths $d<m$,
[…]
@@
$\leq\left(18(c_{0}+c_{1}+9)\log^{2}(T)+28\sqrt{8}\right)4^{m}$
Choosing ${\color[rgb]{0.68,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0,0}\pgfsys@color@cmyk@stroke{0}{0.87}{0.68}{0.32}\pgfsys@color@cmyk@fill{0}{0.87}{0.68}{0.32}c_{A}=18(c_{0}+c_{1}+10)}$ concludes the proof.
@@ D.7 Upper bounding (B)
As discussed in Section 5.3, the main challenge arises in scenarios where a significant shift occurs (as defined in Definition 1), causing the identity of the last safe arm $x_{i}^{\sharp}$ (and consequently the bin $B_{t,m}^{\sharp}$) to change. In order to avoid incurring large cumulative regret before this shift is detected, the algorithm must initiate a replay at the appropriate depth to identify the change in environment. We claim that ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(B)}$ is also of order $4^{m}$ almost surely.
@@
To prove Proposition 8, we analyse the cumulative regret incurred by $B^{\rm{last}}_{l,m}$ with respect to $B_{l,m}^{\sharp}$ by decomposing the rounds of the blocks into bad segments, that is, time intervals during which $B^{\rm{last}}_{l,m}$ incurs significant regret relative to the safe bin $B_{t,m}^{\sharp}$. Importantly, by construction, $B^{\rm{last}}_{l,m}$ is never evicted during the block $[\tau_{l,m},\tau_{l,m+1}-1]$, and thus remains active at depth $m$ throughout. However, since $B^{\rm{last}}_{l,m}$ is only revealed at the end of the block, we must define bad segments relative to all possible bins at depth $m$.
@@
Let $[\tau_{l,m},\tau_{l,m+1}[$ a block, and let $[\tau_{i},\tau_{i+1}[$ be any phase intersecting this block. For all bin $B\in\mathcal{T}_{m}$, define rounds $(s^{m}_{i,j}(B))_{j}$ with $s^{m}_{i,j}(B)\in[\tau_{l,m}\vee\tau_{i},\tau_{i+1}\wedge\tau_{l,m}+8^{m}[$ recursively as follows:
@@
  • $s^{m}_{i,j}(B)$ is the smallest round in $]s^{m}_{i,j-1}(B),\tau_{i+1}\wedge\tau_{l,m+1}[$ such that
    $\sum_{t=s^{m}_{i,j-1}(B)}^{s^{m}_{i,j}(B)}\delta_{t}(B_{t,m}^{\sharp},B)\geq c_{5}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j}(B)-s^{m}_{i,j-1}(B)\right)}\,,$
where $c_{5}$ is a fixed positive numerical constant.
For all bin $B\in\mathcal{T}_{m}$, we define the midpoint of a bad segment $[s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)[$ as the round
@@
Let $B\in\mathcal{T}_{m}$ be a bin and let $[s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)[$ be a bad segment. Then, on $\mathcal{E}_{1}$, any replay starting at round $t_{\mathrm{start}}\in[s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)[$ never evicts $\mathrm{Parent}(B_{t,m}^{\sharp},d)$ for $d\in\bigcap_{t=t_{\mathrm{start}}}^{s^{m}_{i,j+1}(B)}\mathcal{D}_{t}$. Moreover, if a replays at depth $d_{i,j}$ satisfying
@@
Proof of Proposition 9.
Assume a replay evicts $\mathrm{Parent}(B_{t,m}^{\sharp},d)$ with $d\in\bigcap_{t=t_{\mathrm{start}}}^{s^{m}_{i,j+1}(B)}\mathcal{D}_{t}$ and $t_{\mathrm{start}}\geq s_{i,j}^{m}(B)$ before round $s^{m}_{i,j+1}(B)$. Then, by Proposition 1, it means that, on event $\mathcal{E}_{1}$, every arm $x\in B_{t,m}^{\sharp}$ incurs significant regret (Definition 1). This contradicts the definition of $B^{\sharp}_{l,m}$. Indeed, by definition, $x_{t}^{\sharp}$ does not change on $[t_{\mathrm{start}},s_{i,j+1}^{m}]\subset[\tau_{i},\tau_{i+1}]$, so neither does $B_{t,m}^{\sharp}$. Therefore, such replay never evicts $\mathrm{Parent}(B_{t,m}^{\sharp},d)$.
@@
By definition of a bad segment (Definition 12),
$\sum_{s^{m}_{i,j}(B)}^{s^{m}_{i,j+1}(B)}\delta_{t}(B_{t,m}^{\sharp},B)\geq c_{5}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B)\right)}\,.$
@@
$\sum_{s^{m}_{i,j}(B)}^{\tilde{s}^{m}_{i,j}(B)}\delta_{t}(B_{t,m}^{\sharp},B)\leq c_{5}\log(T)\sqrt{2^{m}\left(\tilde{s}^{m}_{i,j}(B)-s^{m}_{i,j}(B)\right)}\,.$
@@
$\sum_{t=\tilde{s}^{m}_{i,j}(B)}^{s^{m}_{i,j+1}(B)}\delta_{t}(B_{t,m}^{\sharp},B)$ $\geq c_{5}\log(T)\sqrt{2^{m}}\left(\sqrt{s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B)}-\sqrt{\tilde{s}^{m}_{i,j}(B)-s^{m}_{i,j}(B)}\right)$
$\geq\frac{c_{5}}{4}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B)\right)}\,,$
@@
Since $B^{\sharp}_{l,m}$ is active during the whole of the bad segment, on $\mathcal{E}_{1}$, we have by Corollary 2, on event $\mathcal{E}_{1}$,
@@
$\geq\left(\frac{c_{5}}{4}-c_{1}\right)\log(T)\sqrt{\left(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B)\right)2^{m}}-\frac{4(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B))}{2^{m}}\,.$
@@
$\sum_{t=\tilde{s}^{m}_{i,j}(B)}^{s^{m}_{i,j+1}(B)}\hat{\delta}_{t}(B_{t,m}^{\sharp},B)\geq\left(\frac{c_{5}}{4}-c_{1}-\frac{32}{\log(T)}\right)\log(T)\sqrt{\left(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B)\right)2^{m}}\,.$
Assuming horizon large enough, e.g. $\log(T)\geq 1$, setting ${\color[rgb]{0.68,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0,0}\pgfsys@color@cmyk@stroke{0}{0.87}{0.68}{0.32}\pgfsys@color@cmyk@fill{0}{0.87}{0.68}{0.32}c_{5}=140+4c_{1}}$ implies
@@
For all block $[\tau_{l,m},\tau_{l,m+1}[\,$, for all bin $B\in\mathcal{T}_{m}$, we define the bad round with respect to $B$ as
$s_{l,m}(B)=$ $\inf\left\{s>\tau_{l,m}\,:\,\sum_{(i,j)\in\mathcal{P}(B,s)}\sqrt{2^{m}(s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B))}>c_{6}\log(T)\sqrt{2^{m}(s-\tau_{l,m})}\right\}$
@@
where $c_{6}$ is a fixed positive numerical constant, and we define the set $\mathcal{P}(B,s)$ as
$\mathcal{P}(B,s)=\big{\{}(i,j)\,:\,i,j\in\mathbb{N}\,,\text{ suchthat }[s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)]\text{ is a bad segment and }s^{m}_{i,j+1}(B)<s\big{\}}\,.$
Now, let $[\tau_{l,m},\tau_{l,m+1})$, be a block, and let $B\in\mathcal{T}_{m}$. We distinguish two cases for this bin.
@@
$\sum_{(i,j)\in\mathcal{P}(B,s_{l,m}(B))}\sqrt{2^{m}(s_{i,j+1}^{m}(B)-s_{i,j}^{m}(B))}$ $\leq c_{6}\log(T)\sqrt{8^{m}2^{m}}$
$\leq c_{6}\log(T)4^{m}\,.$
@@
Let $[\tau_{l,m},\tau_{l,m+1}[$ be a block. We define $\mathcal{E}_{2}(\tau_{l,m})$ as the following event:
$\forall B\in\mathcal{T}_{m}\text{ such that }s_{l,m}(B)<\tau_{l,m+1},\text{ there exists a bad segment }[s_{i,j}^{m}(B),s_{i,j+1}^{m}(B)]\text{such that }$

### v2

@@ D.6 Upper bounding (A)
Our approach begins by upper bounding (A) through a decomposition at depth $d_{0}(t)$ and leverage the properties of our sampling scheme (Algorithm˜2) . In particular, it leverages the fact that at each round $t$, we choose uniformly an active bin at depth $d_{0}(t)$.
@@
where in the last inequality we used the fact that $B_{t,m}\subseteq\mathrm{Parent}(B_{t,m},d_{0}(t))$, $B^{\rm{last}}_{l,m}\subseteq\mathrm{Parent}(B^{\rm{last}}_{l,m},d_{0}(t))$ and Assumption˜1. Note that $\mathrm{Parent}(B_{t,m},d_{0}(t))$ is necessarily active at depth $d_{0}(t)$. Summing this inequality over the block $\tau_{l,m},\tau_{l,m+1}[$,
@@
We are now ready to prove Proposition˜7. The rest of this subsection is dedicated to this purpose.
Thanks to Lemma˜2, it now suffices to bound both the parent’s regret contribution at depth $d_{0}(t)$
@@
Upper bounding (A.1.). Considering all possible value of $d_{0}(t)$, (A.1) can be rewritten using our ordering introduced in Definition˜5: denoting $B_{d}^{(1)}$, $B_{d}^{(1)}\dots,B_{d}^{(2^{d})}$ the bins at depth $d$,
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(A.1)}$
@@
For all bin $B$, $M(s,d,B)$ denotes the last round where bin $B$ is active within a replay at depth $d$ starting at round $s$, $M(s,d,B)\in[s,s+8^{d}]$. If bin $B$ is not evicted during this replay, we define $M(s,d,B)=s+8^{d}$. By abuse of notations, we define $M(\tau_{l,m},m,B)$ as the last round where bin $B$ is retained in $\mathcal{B}_{\mathrm{MASTER}}(m)$, and define $M(\tau_{l,m},m,B)=\tau_{l,m}+8^{m}-1=\tau_{l,m+1}-1$ if it is not evicted during the block $\tau_{l,m},\tau_{l,m+1}[$.
@@
Without any loss of generality, we assume that $\{B_{m}^{(1)},\cdots,B_{m}^{(2^{m})}\}$ is the ordering according to the eviction time of these bins in the block $\tau_{l,m},\tau_{l,m+1}[$ (otherwise we can always re-index the bins), that is,
@@
Moreover, since $B^{\rm{last}}_{l,m}$ and $B_{m}^{(k)}$ are by definition not evicted on interval $\tau_{l,m},M(\tau_{l,m},m,B_{m}^{(k)})]$, we have by Proposition˜5 on event $\mathcal{E}_{1}$
@@
$\leq(c_{0}+c_{1}+8)\log(T)4^{m}$
$\leq c_{2}\log(T)4^{m}\,$,
assuming again horizon $T$ satisfies $\log(T)\geq 1$ and setting $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c_{2}=c_{0}+c_{1}+8$}. Therefore, the regret contribution of replay at depth $m$ is upper bounded as
$\sum_{k=1}^{2^{m}}\sum_{t=\tau_{l,m}}^{M(\tau_{l,m},m,B_{m}^{(k)})}\mathds{1}\!\left\{d_{0}(t)=m\right\}\frac{\delta_{t}\left(B^{\rm{last}}_{l,m},B_{m}^{(k)}\right)}{|\mathcal{B}_{t}(m)|}$ $\leq c_{2}\log(T)4^{m}\sum_{k=1}^{2^{m}}\frac{1}{2^{m}+1-k}$
@@
$\leq(c_{2}+1)\log^{2}(T)4^{m}\,.$ (14)
Regret contribution of replays at depth $d<m$. Fix a depth $d\in[\![0,m-1]\!]$ and a bin $B_{d}^{(k)}$ at this depth. We consider the set of rounds $t>\tau_{l,m}$ for which $d_{0}(t)=d$ and $B_{d}^{(k)}\in\mathcal{B}_{t}(d)$, and partition this set into contiguous intervals (represented as blue intervals in Figure˜2).
@@
Each interval concludes either when $B_{d}^{(k)}$ is evicted from the active set $\mathcal{B}_{t}(d)$, or when the replay itself ends at time $s+8^{d}$. Therefore, each interval spans the range $s,M(s,d,B_{d}^{(k)})]$, where $M(s,d,B_{d}^{(k)})$ is the final round during which $B_{d}^{(k)}$ is active within that specific replay (see Definition˜11).
@@
Following these remarks, we can upper bound the second sum of Section˜D.6 on event $\mathcal{E}_{1}$ as
@@
Now, since the bin $B_{d}^{(k)}$ is active over the interval $s,M(s,d,B_{d}^{(k)})]$, we can apply Proposition˜5 to obtain the desired bound on the event $\mathcal{E}_{1}$.
@@
$\leq(c_{0}+c_{1}+8)\log(T)4^{d}$
$\leq c_{2}\log(T)4^{d}\,$,
assuming horizon $T$ satisfies $\log(T)\geq 1$, and where the second inequality follows from the fact that the length of a replay at depth $d$ stating at round $s$ is at most $8^{d}$. We used the same definition of $c_{2}$ as above. Therefore, we have
@@
$\leq c_{2}\log(T)\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{d=0}^{m-1}\sum_{k=1}^{2^{d}}\frac{1}{2^{d}-k+1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}4^{d}\,\bigg|\,\mathcal{F}_{\tau_{l,m}}\right]$
$\leq c_{2}\log(T)\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{d=0}^{m-1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}(\log(2^{d})+1)4^{d}\,\bigg|\,\mathcal{F}_{\tau_{l,m}}\right]$
$\leq c_{2}\log^{2}(T)\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{d=0}^{m-1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}4^{d}\,\bigg|\,\mathcal{F}_{\tau_{l,m}}\right]$
$\leq c_{2}\log^{2}(T)\mathbb{E}\left[\sum_{d=0}^{m-1}\sum_{s=\tau_{l,m}}^{\tau_{l,m+1}-1}R_{s,d}4^{d}\,\bigg|\,\mathcal{F}_{\tau_{l,m}}\right]\,$,
@@
We now leverage the fact that relatively few replays occur in expectation, and the fact that the Bernoulli random variables $R_{s,d}$ are sampled independently of the observations collected during the episode. More precisely, all Bernoulli variables are drawn i.i.d. at the beginning of the block, i.e. at round $t=\tau_{l,m}$. This allows us to treat the scheduling decisions as fixed ahead of time. Furthermore, by the design of the replay scheduling mechanism in Algorithm˜2,
@@
$\leq 6\sqrt{8}c_{2}\log^{2}(T)4^{m}\,.$ (16)
From Section˜D.6, combining Sections˜D.6 and D.6 gives the desired rate for (A.1),
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(A.1)}\leq 18(c_{2}+1)\log^{2}(T)4^{m}\,.$ (17)
Upper bounding (A.2). To bound this term, we apply the same reasoning previously used to control (A.1), and in particular the partition argument of Figure˜2. By observing that each replay at depth $d_{0}(t)=d$ is initiated by a round $s$ such that $R_{s,d}=1$, and by using the fact that each replay at depth $d$ lasts at most $8^{d}$ rounds, we distinguish the cumulative discretization bias incurred at depth $m$ and the one incurred at other depths $d<m$,
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(A.2)}$ $=\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}\sum_{t=\tau_{l,m}}^{\tau_{l,m+1}-1}\frac{4}{2^{d_{0}(t)}}\,\bigg|\,\mathcal{F}_{\tau_{l,m}}\right]$
@@
$\leq\left(18(c_{2}+1)\log^{2}(T)+28\sqrt{8}\right)4^{m}$
Choosing $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c_{A}=18(c_{2}+2)$} concludes the proof.
@@ D.7 Upper bounding (B)
As discussed in Section˜5.3, the main challenge arises in scenarios where a significant shift occurs (as defined in Definition˜1), causing the identity of the last safe arm $x_{i}^{\sharp}$ (and consequently the bin $B_{t,m}^{\sharp}$) to change. In order to avoid incurring large cumulative regret before this shift is detected, the algorithm must initiate a replay at the appropriate depth to identify the change in environment. We claim that $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(B)$} is also of order $4^{m}$ almost surely.
@@
To prove Proposition˜8, we analyse the cumulative regret incurred by $B^{\rm{last}}_{l,m}$ with respect to $B_{l,m}^{\sharp}$ by decomposing the rounds of the blocks into bad segments, that is, time intervals during which $B^{\rm{last}}_{l,m}$ incurs significant regret relative to the safe bin $B_{t,m}^{\sharp}$. Importantly, by construction, $B^{\rm{last}}_{l,m}$ is never evicted during the block $\tau_{l,m},\tau_{l,m+1}-1]$, and thus remains active at depth $m$ throughout. However, since $B^{\rm{last}}_{l,m}$ is only revealed at the end of the block, we must define bad segments relative to all possible bins at depth $m$.
@@
Let $\tau_{l,m},\tau_{l,m+1}[$ a block, and let $\tau_{i},\tau_{i+1}[$ be any phase intersecting this block. For all bin $B\in\mathcal{T}_{m}$, define rounds $(s^{m}_{i,j}(B))_{j}$ with $s^{m}_{i,j}(B)\in[\tau_{l,m}\vee\tau_{i},\tau_{i+1}\wedge\tau_{l,m}+8^{m}[$ recursively as follows:
@@
  • $s^{m}_{i,j}(B)$ is the smallest round in $s^{m}_{i,j-1}(B),\tau_{i+1}\wedge\tau_{l,m+1}[$ such that
    $\sum_{t=s^{m}_{i,j-1}(B)}^{s^{m}_{i,j}(B)}\delta_{t}(B_{t,m}^{\sharp},B)\geq c_{3}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j}(B)-s^{m}_{i,j-1}(B)\right)}\,$,
where $c_{3}$ is a fixed positive numerical constant.
For all bin $B\in\mathcal{T}_{m}$, we define the midpoint of a bad segment $s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)[$ as the round
@@
Let $B\in\mathcal{T}_{m}$ be a bin and let $s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)[$ be a bad segment. Then, on $\mathcal{E}_{1}$, any replay starting at round $t_{\mathrm{start}}\in[s^{m}_{i,j}(B),s^{m}_{i,j+1}(B)[$ never evicts $\mathrm{Parent}(B_{t,m}^{\sharp},d)$ for $d\in\bigcap_{t=t_{\mathrm{start}}}^{s^{m}_{i,j+1}(B)}\mathcal{D}_{t}$. Moreover, if a replays at depth $d_{i,j}$ satisfying
@@
Proof of Proposition˜9.
Assume a replay evicts $\mathrm{Parent}(B_{t,m}^{\sharp},d)$ with $d\in\bigcap_{t=t_{\mathrm{start}}}^{s^{m}_{i,j+1}(B)}\mathcal{D}_{t}$ and $t_{\mathrm{start}}\geq s_{i,j}^{m}(B)$ before round $s^{m}_{i,j+1}(B)$. Then, by Proposition˜1, it means that, on event $\mathcal{E}_{1}$, every arm $x\in B_{t,m}^{\sharp}$ incurs significant regret (Definition˜1). This contradicts the definition of $B^{\sharp}_{l,m}$. Indeed, by definition, $x_{t}^{\sharp}$ does not change on $t_{\mathrm{start}},s_{i,j+1}^{m}]\subset[\tau_{i},\tau_{i+1}]$, so neither does $B_{t,m}^{\sharp}$. Therefore, such replay never evicts $\mathrm{Parent}(B_{t,m}^{\sharp},d)$.
@@
By definition of a bad segment (Definition˜12),
$\sum_{s^{m}_{i,j}(B)}^{s^{m}_{i,j+1}(B)}\delta_{t}(B_{t,m}^{\sharp},B)\geq c_{3}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B)\right)}\,.$
@@
$\sum_{s^{m}_{i,j}(B)}^{\tilde{s}^{m}_{i,j}(B)}\delta_{t}(B_{t,m}^{\sharp},B)\leq c_{3}\log(T)\sqrt{2^{m}\left(\tilde{s}^{m}_{i,j}(B)-s^{m}_{i,j}(B)\right)}\,.$
@@
$\sum_{t=\tilde{s}^{m}_{i,j}(B)}^{s^{m}_{i,j+1}(B)}\delta_{t}(B_{t,m}^{\sharp},B)$ $\geq c_{3}\log(T)\sqrt{2^{m}}\left(\sqrt{s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B)}-\sqrt{\tilde{s}^{m}_{i,j}(B)-s^{m}_{i,j}(B)}\right)$
$\geq\frac{c_{3}}{4}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B)\right)}\,$,
@@
Since $B^{\sharp}_{l,m}$ is active during the whole of the bad segment, on $\mathcal{E}_{1}$, we have by Corollary˜2, on event $\mathcal{E}_{1}$,
@@
$\geq\left(\frac{c_{3}}{4}-c_{1}\right)\log(T)\sqrt{\left(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B)\right)2^{m}}-\frac{4(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B))}{2^{m}}\,.$
@@
$\sum_{t=\tilde{s}^{m}_{i,j}(B)}^{s^{m}_{i,j+1}(B)}\hat{\delta}_{t}(B_{t,m}^{\sharp},B)\geq\left(\frac{c_{3}}{4}-c_{1}-\frac{32}{\log(T)}\right)\log(T)\sqrt{\left(s^{m}_{i,j+1}(B)-\tilde{s}^{m}_{i,j}(B)\right)2^{m}}\,.$
Assuming horizon large enough, e.g. $\log(T)\geq 1$, setting $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c_{3}=140+4c_{1}$} implies
@@
For all block $\tau_{l,m},\tau_{l,m+1}[\,$, for all bin $B\in\mathcal{T}_{m}$, we define the bad round with respect to $B$ as
$s_{l,m}(B)=$ $\inf\left\{s>\tau_{l,m}\,:\,\sum_{(i,j)\in\mathcal{P}(B,s)}\sqrt{2^{m}(s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B))}>c_{4}\log(T)\sqrt{2^{m}(s-\tau_{l,m})}\right\}$
@@
where $c_{4}$ is a fixed positive numerical constant, and we define the set $\mathcal{P}(B,s)$ as
[…]
Now, let $\tau_{l,m},\tau_{l,m+1})$, be a block, and let $B\in\mathcal{T}_{m}$. We distinguish two cases for this bin.
@@
$\sum_{(i,j)\in\mathcal{P}(B,s_{l,m}(B))}\sqrt{2^{m}(s_{i,j+1}^{m}(B)-s_{i,j}^{m}(B))}$ $\leq c_{4}\log(T)\sqrt{8^{m}2^{m}}=c_{4}\log(T)4^{m}\,.$
@@
Let $\tau_{l,m},\tau_{l,m+1}[$ be a block. We define $\mathcal{E}_{2}(\tau_{l,m})$ as the following event:
[…]